# Enhancing variational quantum algorithms by balancing training on classical and quantum hardware

### v1

@@ II.3.1 Parameter-shift Rule
where we defined $U_{<k}=U_{1}W_{1}U_{2}W_{2}\ldots U_{k-1}W_{k-1}$ and $U_{\geq k}=U_{k}W_{k}U_{k+1}W_{k+1}\ldots U_{n}W_{n}$. PSR can be employed to obtain a closed-form expression of Eq.(II.3.1) when the eigenspectrum of the generator $G_{k}$ either has only two unique eigenvalues or is symmetric and evenly spaced [71, 28, 72, 73, 30].
@@
where we have defined $|\psi\rangle=U_{\geq k}|0\rangle$ and $Q={U}^{\dagger}_{<k}{\hat{O}}U_{<k}$. Finally using Eq.(6) and substituting $\theta=\pi/4r$, the partial derivative simplifies to
$\frac{\partial{\ell}}{\partial{\theta_{k}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{k})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{k})\Big{)},$ (8)
a difference between the loss functions evaluated on either side of the parameter with a specific eigenvalue-related shift.
By measuring the loss function at two shifted values of the parameter $\theta_{k}$, the partial derivative can be estimated accurately upto shot noise. For more general eigenspectrum with equispaced eigenvalues, the parameter shift-rule can be generalized as shown in [30, 73]. Markovich et.al. have further extended this procedure to irregeular eigenspectrum as well [74].
The computation of a gradient for a single parameter in a quantum circuit requires executing the circuit at least twice, apart from the measurement overhead to obtain expectation value accurately. In quantum machine learning (QML), backpropagation is not directly applicable because it relies on the ability to store or reuse information [33], which poses scalability challenges. This limitation is illustrated through a specific example. Assuming each circuit iteration takes 1.48 milliseconds [75], equivalent to the latest coherence time of superconducting qubits, one iteration for a circuit with a billion parameters would take approximately $2\times 10^{9}\times 1.48$ ms, or 296 million seconds (3 days , 10 hours and 19 minutes). This demonstrates the poor scalability of PSR for large-scale systems. For certain examples of PQC construction, the training may scale more favorably. One such example is proposed in Bowles et. al to address the scalability issue by introducing structured circuits consisting of blocks of commuting parametrized quantum gates [76]. However, the utility of such structured PQCs is still not fully understood.
II.3.2 $\mathfrak{g}$-sim Method
In this section, we briefly explain the details of $\mathfrak{g}$-sim method proposed by Goh et al. [31], that is fully classical in nature. $\mathfrak{g}$-sim method is based on the study of the Lie group and the associated Lie algebra $g$, which is generated by the parametrized quantum circuit. Note that the foundational techniques used in [31] for this gradient estimation method are well-established by Somma et al. [77], Goh et al. primary contributed to reframe these techniques in a modern context tailored to the quantum computing community [31].
Figure 2: Quantum circuit for $\mathfrak{g}$-sim algorithm corresponding to the unitary in Eq. (9)
Consider a problem in which an $n$-qubit state $\rho$, acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ is sent through a PQC $U(\boldsymbol{\theta})$ of the form
where $\boldsymbol{\theta}=[\theta_{11},\theta_{12},\ldots,\theta_{LK}]$ is a set of trainable real-valued parameters, and $\{H_{1},H_{2},\ldots,H_{K}\}$ are $K$ Hermitian operators that are the gate generators of the circuit, as shown in Fig 2.
$\mathcal{G}=\{e^{iA}:iA\in\mathfrak{g}\},$
$\mathfrak{g}$-sim method is applicable only under two assumptions [31]:
  • either the measurement operator or the initial state of a variational quantum circuit belong to the DLA of the ansatz
  • the dimension of DLA scales polynomially with the number of qubits.
@@ II.3.2 $\mathfrak{g}$-sim Method
Input: $T_{max},\alpha,H=\sum_{i}^{g_{max}}a_{i}G_{i},\linebreak U(\vec{\theta})=e^{-iG_{1}\theta_{1}}\ldots e^{-iG_{n}\theta_{n}}$
@@
1                         $H_{k}^{m-1}=e^{iG_{m}\theta_{m}}(a_{k}^{m}G_{k})e^{-iG_{m}\theta_{m}}$
                         $\quad\quad=\sum_{j}b^{m-1}_{j}G_{j}$
@@ 7             Obtain gradients $\gradient_{\vec{\theta}}C$ classically using an automatic differentiation framework
8             $\vec{\theta}\leftarrow\vec{\theta}-\alpha\gradient_{\vec{\theta}}C$
@@ 10            $i\leftarrow i+1$
Using the above DLA structure we demonstrate an implementation of the full $\mathfrak{g}$-sim sub-routine in action for VQE in Alg.1. We choose a Hamiltonian with a DLA of polynomial dimension. The ansatz is chosen as $U(\boldsymbol{\theta})=e^{-i\theta_{1}G_{1}}\ldots e^{-i\theta_{n}G_{n}}$, where the generators are from the same DLA, allowing us to efficiently use the DLA structure in our computation.
The crucial factor facilitating the use of DLA structure is the Baker-Campbell-Hausdorff (BCH) formula [79] in Step. 1, which is given by
@@
The limitations of the PSR and the efficiency of $\mathfrak{g}$-sim provide motivation for combining them into a hybrid approach to achieve efficient gradient evaluation. Apart from scalability and resource challenges in gradient evaluation, the training of quantum models remains challenging, primarily due to the vanishing gradient phenomenon, also known as the BP problem, where gradients decay exponentially as the number of qubits increases, making larger models effectively untrainable. We expand on this point in the following subsection.
II.4 Barren Plateau
Optimization in quantum algorithms encounters significant challenges due to the intrinsic limitations imposed by quantum measurements. Classical optimization methods rely on precise loss evaluations or accurate gradient information, both of which are compromised by the finite precision and stochastic nature of quantum measurements.
@@ II.4 Barren Plateau
Barren plateaus (BPs) in quantum variational algorithms are characterized by regions in the parameter space where the gradient of the loss function becomes exponentially small, severely impeding the efficient optimization of quantum circuits [81, 82, 83]. The presence of a BP landscape implies that, for most parameter configurations, an exponentially large number of measurement samples is required to reliably estimate the loss gradient, which is essential for effective optimization.
@@
The Variational Quantum Algorithm (VQA) is believed to exhibit a BP when variations in the model parameters $\theta$ in the unitary operation $U(\theta)$ lead to only exponentially small changes in the loss function $l_{\theta}(\rho,\hat{O})$, or magnitude of gradients $\partial l_{\theta}(\rho,\hat{O})/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\theta$) [34, 82, 83]. Cerezo et al., in their investigation of cost function locality, established that the inclusion of global observables $(\hat{O})$ in the cost function results in the emergence of exponentially vanishing gradients, irrespective of the circuit’s shallow depth [84]. Additional factors contributing to this phenomenon include the degree of entanglement [85, 86] and the presence of noise [87]. An unified theoretical framework, employing group theoretical principles, covering above factors were presented in [82, 83].
@@
Beyond theoretical insights, substantial efforts have been devoted to developing practical strategies for mitigating barren plateaus [88]. These include advanced initialization techniques, such as layer-wise training [89], Bayesian-inspired parameter settings [90], meta-learning for initializing parameters [91], initializing circuits with sequences of identity blocks [92], and transfer-learning-based initialization approaches [93]. Alternative optimization frameworks have also been introduced, including adaptive learning rate strategies [94] and novel line search methods designed to facilitate navigation through flat loss landscapes [95]. Mhiri et.al. provided an unified analysis on warm-starting approaches in BP-affected loss landscape [96]. Classical neural network-based methodologies have also been explored, such as generating parameters for parameterized quantum circuits (PQCs) using neural networks [97, 98].
Empirical studies, such as those conducted in Ref. [99], have validated the effectiveness of these approaches in reducing the prevalence of barren plateaus in practical applications, including quantum neural networks. Additionally, targeted strategies for specific ansatz designs, such as Hamiltonian variational ansatz and parameterized matchgate circuits, have demonstrated their potential in mitigating or entirely avoiding barren plateaus [100, 101].
@@
However, recent research has shed light on the classical simulability of BP-free models, offering valuable insights into their underlying structure. Cerezo et al. highlights that BP-free models are classically simulable when classical data is gathered during an initial data acquisition phase from the quantum device. This conclusion stems from the fact that barren plateaus arise from the curse of dimensionality, with mitigation strategies effectively reducing the problem to subspaces that are classically tractable [38].
Finally, using classical processing to improve the results from quantum device is not unique to our work, and has been used in [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112]. In fact, a recent paper also deals with the idea of processing different parts of the quantum circuit using classical and quantum hardware with Clifford Perturbation Theory [113]. However, our approach of combining $\mathfrak{g}$-sim and the PSR in a hybrid iterative fashion offers a balanced method to address the barren plateau problem, improve resource efficiency, and maintain non-classical simulability. This makes it a promising area for further research and development in quantum computing.
III Our Contributions
Gathering insights from current literature, we find extensive challenges across all building blocks of a variational quantum algorithm trained on today’s hardware. In this paper, we address issues related to choice of PQC, resource-inefficiency of gradient evaluation and trainability of large models in regards to scaling the number of parameters and the concurrent barren plateau issues that appear. We go through each of the improvements in a sequential manner highlighting clear instances of where and how the improvements are obtained using detailed numerical analysis.
III.1 PQC Selection
Our choice of PQC is motivated from the standpoint of reducing quantum resources for training as well as avoiding Barren Plateaus. In general, simulating the full density matrix requires keeping track of all the $n$-qubit Pauli operators which scales as $14^{n}-$ 1. However, to effectively utilize $\mathfrak{g}$-sim without incurring exponential computational and memory overhead, it is essential to restrict the training ansatz and measurement operator associated with a QML task to a poly-DLA. This can be a challenge, for example in VQE if the ground state lies far outside the polynomial DLA(poly-DLA) being explored (as shown in Table 3 in Appendix A), or a QML task whose accuracy gets limited by the polynomial search space (as shown in Fig.1).
Figure 3: The HELIA comprises two blocks of quantum gates: the orange block or Block Q ($U_{q}$) consists of layers of HEA (non-DLA part), followed by the blue block or Block G ($U_{g}$), which includes multi-qubit Pauli rotation gates with generators forming the basis of a task-specific DLA. The final measurement operator also belongs to the same DLA
To elevate $\mathfrak{g}$-sim to a higher expressibility, we propose a hybrid ansatz consisting of a parameterized non-DLA block $U_{q}$, which cannot be classically simulated, followed by a parameterized $\mathfrak{g}$-sim block $U_{g}$ that can be classically simulated within the $\mathfrak{g}$-sim framework. A specific ansatz falling within this proposal is the Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), shown in Fig. 3, where $U_{q}$ represents the hardware efficient ansatz (HEA). $U_{g}$ is composed of multi-qubit Pauli rotations $\exp{-i\theta P}$ ), where the Paulis $P$ form an orthonormal basis of a chosen poly-DLA. Both HEA [62] and Lie algebra Supported Ansatz [83] are known in literature, however our proposal aims at combining them into a single ansatz, which we will show has several important advantages.
This specific ordering of gates is important for our protocol, as we require the gates and measurement operator chosen from the poly-DLA to be adjacent in the quantum circuit for $\mathfrak{g}$-sim to be used efficiently. If the order is reversed, the Heisenberg-evolved operator obtained after applying the HEA gates on the measurement operator will have terms from outside the poly-DLA, making $\mathfrak{g}$-sim quite ineffective. In our examples of VQE and Classification, we use a YZ linear ansatz, shown in Fig. 1 (which belongs to the HEA class), with varying layers, and the DLA is motivated by the problem being solved.
Figure 4: VQE performed for a 6-qubit TFIM Hamiltonian. Only using HEA using PSR (shown in orange) or multi-qubit Pauli rotation gates using $\mathfrak{g}$-sim (shown in blue) is not enough to reach exact ground state (shown in red). But combining both methods into an Alternate training procedure (in green) allows us to reach the correct solution.
The primary advantage of HELIA lies in its ability to use PSR for computing gradients in the orange block of gates, while utilizing $\mathfrak{g}$-sim for the blue block in Fig 3. Additionally, we introduce two novel training schemes and demonstrate their superior performance compared to standard PSR or $\mathfrak{g}$-sim independently. This improvement is illustrated in Fig.4 for a 6-qubit Transverse Field Ising Model (TFIM) Hamiltonian, where using only the orange block (HEA via PSR) or the blue block (multi-qubit Pauli gates via $\mathfrak{g}$-sim ) fails to reach the target energy in 500 iterations. However, combining the two blocks using an Alternate training scheme (discussed in Sec.III.2) successfully converges to the exact ground state energy. Finally, we investigate the barren plateau phenomenon and empirically demonstrate that our ansatz exhibits a slower decay of gradients with increasing qubit count, compared to a deep hardware-efficient circuit, thereby enabling the effective training of larger qubit models.
III.2 Training Scheme
For the selected PQC, two distinct training schemes are employed: Alternate and Simultaneous. We evaluate the proposed protocols on two different VQA applications: ground state estimation using VQE and the Quantum Phase Classification. The core idea in both the schemes is to exploit the ansatz structure and distribute the resources required for gradient evaluation to both classical and quantum hardware. This delegation not only significantly reduces the number of calls to the quantum hardware, but also increases the accuracy as shown in the results of VQE in Table 2 and 3 of Appendix A. In the subsequent text we will interchangeably refer to the orange block of gates as Block Q or PSR Block, and the blue block as Block G or $\mathfrak{g}$-sim Block, based on the protocol used in obtaining the associated gradients.
III.2.1 Alternate Training
This section outlines a training protocol that integrates both the PSR and $\mathfrak{g}$-sim to develop a hybrid algorithm for gradient evaluation and quantum model training. As described in Algorithm 2 and depicted in Fig. 5, this protocol alternates the gradient evaluation process between classical and quantum hardware.
In each optimization step we update the parameters of Block Q (from $\theta_{i}$ to $\theta_{i+1}$) in Fig.3 using standard PSR. This is followed by a step of updating the parameters of Block G (from $\phi_{i}$ to $\phi_{i+1}$) using $\mathfrak{g}$-sim based on the updated parameters $\theta_{i+1}$. The optimization process is repeated iteratively until convergence.
Figure 5: Workflow of Alternate training protocol based on Alg.2
Let’s consider the case where the Block Q is composed of YZ linear ansatz (forming an unitary $U_{q}(\theta_{i})$) with $p$ parameters and the Block G has $g$ parametrized gates corresponding to a DLA of size $g$ (forming an unitary $U_{g}(\phi_{i})$), and the initial state is $|\psi\rangle$. Obtaining the gradients of the Block Q will require atleast $2p$ different runs of the quantum circuit for PSR. Using the gradients, we update Block Q to $U_{q}(\theta_{i+1})$.
1
Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
Initialize: $\mathbf{\phi}_{0},\mathbf{\theta}_{0}\sim Norm(0,1);i=0$
Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
begin
       while $i\leq T_{s}$ do
             PSR Step
3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
             $\mathfrak{g}$-sim Step
5             for $\hat{O}_{m}\in\mathfrak{g}_{DLA}$ do
9             $i\leftarrow i+1$
10
11
12
Algorithm 2 Alternate Training for VQE
To evaluate gradients of the Block G, we measure the expectation value of all $g$ Pauli operators on the state $|\psi(\theta_{i+1})\rangle=U_{q}(\theta_{i+1})|\psi\rangle$ on quantum hardware. This data then feeds into the $\mathfrak{g}$-sim algorithm which can evaluate the gradients entirely on classical hardware using an automatic differentiation framework [69, 70]. Finally, the parameters of Block G are updated from $\phi_{i}$ to $\phi_{i+1}$, finishing one iteration. Hence, each of these iterations, costs us $2p+g$ circuit evaluations (ignoring the factor of number of shots required for estimating the expectation values). This is in contrast with the $2p+2g$ circuit evaluations required per iteration to train the entire circuit via PSR alone.
III.2.2 Simultaneous Training
In this section, we propose a modification to the Alternate training protocol that enables the simultaneous updating of parameters in each iteration.
The pseudo-algorithm is shown in Algorithm 3. The main difference with the Alternate training scheme is, here we update the Block G gates based on $\theta_{i}$ (not the updated $\theta_{i+1}$). This is essentially equivalent to obtaining gradients for the full ansatz by PSR, and updating all parameters simultaneously. However, we significantly reduce the resource requirements by delegating a portion of the gradient estimation task to classical hardware via $\mathfrak{g}$-sim . The required circuit evaluations per iteration is also the same as the above Alternate scheme, as the only difference comes in by using the non-updated unitary $U_{q}(\theta_{i})$ in the measurement step before $\mathfrak{g}$-sim .
1
Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
Initialize: $\mathbf{\phi_{0}},\mathbf{\theta_{0}}\sim Norm(0,1);i=0$
Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
begin
       while $i\leq T_{s}$ do
             PSR Step
3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
             $\mathfrak{g}$-sim Step
5             for $\hat{O}_{m}\in\mathfrak{g}_{DLA}$ do
9             $i\leftarrow i+1$
10
11
12
Algorithm 3 Simultaneous Training for VQE
In practice, we find the Alternate scheme for a fixed number of iterations followed by several iterations of the Simultaneous training provides the most optimal results. We will expand on this point in the subsequent sections.
IV Results
To evaluate the effectiveness of our method we employ several metrics,
Error:
$\text{Error}=\frac{E_{g}^{t}-E_{g}^{*}}{E_{g}^{*}}$ (13)
where we define $E_{g}^{t}$ as the lowest possible energy achieved in the current trial and $E_{g}^{*}$ is an estimate of the true ground state energy of the Hamiltonian. For large Hamiltonians (beyond 14 qubits) it is computationally impractical to numerically diagonalize it and obtain the true ground state energy. Instead, taking inspiration from [31], we use the lowest energy reached across all trials for the corresponding Hamiltonian as an estimate of the true ground state energy $E_{g}^{*}$.

### v2

@@ Appendix A Parameter-shift Rule
where we defined $U_{<k}=U_{1}W_{1}U_{2}W_{2}\ldots U_{k-1}W_{k-1}$ and $U_{\geq k}=U_{k}W_{k}U_{k+1}W_{k+1}\ldots U_{n}W_{n}$. PSR can be employed to obtain a closed-form expression of Eq.(A) when the eigenspectrum of the generator $G_{k}$ either has only two unique eigenvalues or is symmetric and evenly spaced [65, 28, 104, 68, 30].
@@
where we have defined $|\psi\rangle=U_{\geq k}|0\rangle$ and $Q={U}^{\dagger}_{<k}{\hat{O}}U_{<k}$. Finally using Eq.(19) and substituting $\theta=\pi/4r$, the partial derivative simplifies to a difference between the loss functions evaluated on either side of the parameter with a specific eigenvalue-related shift
$\frac{\partial{\ell}}{\partial{\theta_{k}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{k})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{k})\Big{)}.$ (21)
It is important to note that, despite its similarity to the finite difference approximation, the above formula is infact an exact estimation of the partial derivative upto shot noise.
Appendix B $\mathfrak{g}$-sim Method
@@
Input: $T_{max},\eta,H=\sum_{i}^{g_{max}}a_{i}G_{i},\linebreak U(\vec{\theta})=e^{-iG_{1}\theta_{1}}\ldots e^{-iG_{n}\theta_{n}}$
@@
1                         $H_{k}^{m-1}=e^{iG_{m}\theta_{m}}(a_{k}^{m}G_{k})e^{-iG_{m}\theta_{m}}=\sum_{j}b^{m-1}_{j}G_{j}$
@@ 7             Obtain gradients $\gradient_{\vec{\theta}}C$ classically using an automatic differentiation framework
8             $\vec{\theta}\leftarrow\vec{\theta}-\eta\gradient_{\vec{\theta}}C$
@@ 10            $i\leftarrow i+1$
Using the above DLA structure we demonstrate an implementation of the full $\mathfrak{g}$-sim sub-routine in action for VQE in Alg.3. We choose a Hamiltonian with a DLA of polynomial dimension. The ansatz is chosen as $U(\boldsymbol{\theta})=e^{-i\theta_{1}G_{1}}\ldots e^{-i\theta_{n}G_{n}}$, where the generators are from the same DLA, allowing us to efficiently use the DLA structure in our computation.
The crucial factor facilitating the use of DLA structure is the Baker-Campbell-Hausdorff (BCH) formula [105] in Step. 3, which is given by
@@
Appendix C Overview of Barren Plateaus (BPs) and proposed mitigation techniques
@@
BPs are characterized by regions in the parameter space where the gradient of the loss function becomes exponentially small, severely impeding the efficient optimization of quantum circuits [95, 96, 94]. The presence of a BP landscape implies that, for most parameter configurations, an exponentially large number of measurement samples is required to reliably estimate the loss gradient, which is essential for effective optimization.
@@
VQA is believed to exhibit a BP when variations in the model parameters $\theta$ lead to only exponentially small changes in the loss function $l_{\theta}(\rho,\hat{O})$, or magnitude of gradients $\partial l_{\theta}(\rho,\hat{O})/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\theta$[31, 96, 94]. Cerezo et al., in their investigation of cost function locality, established that the inclusion of global observables $(\hat{O})$ in the cost function results in the emergence of exponentially vanishing gradients, irrespective of the circuit’s shallow depth [107]. Additional factors contributing to this phenomenon include the degree of entanglement [108, 109, 110] and the presence of noise [111]. A unified theoretical framework, employing group theoretical principles, covering above factors were presented in Refs. [96, 94].
@@
Beyond theoretical insights, substantial efforts have been devoted to developing practical strategies for mitigating BPs [112]. These include advanced initialization techniques, such as layer-wise training [113], Bayesian-inspired parameter settings [114], meta-learning for initializing parameters [115], initializing circuits with sequences of identity blocks [116], and transfer-learning-based initialization approaches [117]. Alternative optimization frameworks have also been introduced, including adaptive learning rate strategies [118] and novel line search methods designed to facilitate navigation through flat loss landscapes [119]. Mhiri et.al. provided an unified analysis on warm-starting approaches in BP-affected loss landscape [120]. Classical neural network-based methodologies have also been explored, such as generating parameters for PQCs using neural networks [121, 122].
Empirical studies, such as those conducted in Ref. [123], have validated the effectiveness of these approaches in reducing the prevalence of BPs in practical applications, including QNNs. Enrique et al. have shown that the optimization landscapes of certain quantum circuit architectures inspired by classical tensor network structures such as, tree tensor networks (qTTN) and the multiscale entanglement renormalization ansatz (qMERA) are free of BPs [110]. Furthermore, the HVA and parameterized matchgate circuits have also been shown to effectively mitigate or entirely circumvent BPs [101, 124].
@@
Recent research has shed light on the classical simulability of BP-free models, offering valuable insights into their underlying structure. Cerezo et al. suggested that BP-free models might be classically simulable when classical data is gathered during an initial data acquisition phase from the quantum device [35]. This hypothesis stems from the observation that BPs arise from the curse of dimensionality, with mitigation strategies effectively reducing the problem to subspaces that are classically tractable.
Appendix D Numerical Experiments on XY Hamiltonian
The success, relative error and QPU calls reduction have been plotted for the HELIA configuration with 1 layer of YZ linear ansatz in Block Q in Fig. 5. For the configurations with 3,6 and 9 YZ linear layers we plot the corresponding metrics in Fig.10.
(a) Success, Relative Error and QPU Call Reduction for 3 YZ linear layers in Block Q
(a) Success, Relative Error and QPU Call Reduction for 6 YZ linear layers in Block Q
(a) Success, Relative Error and QPU Call Reduction for 9 YZ linear layers in Block Q